# MergeBench: A Benchmark for Merging Domain-Specialized LLMs

### v2

@@ 1 Introduction
To address the limitations of existing model merging evaluations, we introduce MergeBench, a scalable and comprehensive benchmark designed to rigorously assess merging performance, illustrated in Figure 1. First, MergeBench improves model selection by adopting state-of-the-art, open-source language models as base models. Specifically, we include both pretrained and instruction-tuned versions of Llama-3.2-3B, Llama-3.1-8B [17], Gemma-2-2B, and Gemma-2-9B [62], resulting in a total of eight base models. Second, we construct a more challenging and representative task suite for evaluating merged models. Each base model is further finetuned on one of five carefully selected task categories, including instruction following, mathematics, multilingual understanding, coding and safety, to produce specialized models with minimal skill overlap [^1] All of the 40 specialized models are open-sourced at https://huggingface.co/MergeBench.. By standardizing the finetuning and evaluation procedures, MergeBench ensures a fair and reproducible platform for comparing model merging algorithms. In addition to multi-task performance, MergeBench evaluates retention of pretrained generalization through forgetting analysis and reports runtime efficiency, offering a comprehensive view of both utility and computational cost of existing merging algorithms.
@@ 2 Background
Localize-and-Stitch [22] i) Localization: Train binary mask to identify the most relevant parameters for each task. Dataless localization: When no data available, retain largest top-$k$ parameters in task vectors. ii) Stitch: Only stitch the localized regions back onto the base model.
Methods. We evaluate and compare eight representative model merging methods, summarized in Table 1. The development of model merging techniques begins with the study of how to effectively assign merging coefficients to the constituent models. Model Soup, Task Arithmetic, Fisher Merging and RegMean exemplify this approach. As the field has evolved, researchers have identified sparsity in task vectors as critical to reducing interference during merging. Since finetuning often results in redundant or noisy parameter changes [44], sparse merging techniques aim to suppress uninformative updates. These methods typically involve sparsification alongside merging coefficient tuning. More details about each method is included in Appendix A.
@@ 3 MergeBench
MergeBench provides a framework to evaluate model merging methods with three key designs: task coverage, model selection, and training and evaluation procedure. We define diverse task categories (Section 3.1), build specialized models from open-source LLMs (Section 3.2), and apply standardized training and evaluation strategies (Section 3.3) to ensure fair and reproducible evaluation.
@@ 3.3 Data Construction
Training. Since the base models already go through the pretraining stage, our training primarily focuses on post-training. For most task categories, we apply supervised finetuning (SFT) to align the base models to domain-specific behaviors. To better reflect realistic scenarios where specialized models may be developed asynchronously using different methods, we adopt additional training strategies where appropriate. Specifically, for mathematics on the 8B and 9B models, we further apply Group Relative Policy Optimization (GRPO) [56] on top of SFT to enhance the models’ reasoning capabilities. A summary of the training data and methods is provided in Table 2, with detailed data statistics in Section B.1 and training configurations available in Section C.1.
@@
Merging. To enable fair comparison across merging algorithms, we standardize the evaluation protocol by unifying their data requirements. Except for Model Soup, all other methods depend on auxiliary data, which we categorize into two types. First, some methods require additional training data to compute model-specific statistics or perform optimization. This applies to Fisher Merging, RegMean and Localize-and-Stitch. For these methods, we uniformly sample 1,000 examples from the original training set. Second, several methods require validation data to tune hyperparameters, typically scaling factors that control the contribution of each task vector and sparsity levels that control the proportion of parameters retained. This category includes Task Arithmetic, TIES Merging, DARE, Consensus TA and Dataless Localize-and-Stitch. We tune these hyperparameters based on performance on surrogate validation tasks, with details in Section B.2.
@@
Evaluation. To assess the performance of the merged models, we curate a comprehensive evaluation suite covering all five task categories. The datasets and evaluation metrics used for each category are summarized in Table 3. In addition to task performance, we evaluate the efficiency of each merging algorithm by reporting their wall-clock time, allowing for a holistic comparison that considers not only effectiveness but also the practical cost of applying each method.
@@ 4 Evaluation of Merging Methods
To provide a comprehensive evaluation of merging algorithms, we assess their performance along three key dimensions. First, we measure the multi-task performance of the merged models on the five target tasks, as detailed in Section 4.1. Second, we assess forgetting of base model knowledge, evaluating how merging impacts the model’s generalization beyond the specialized tasks in Section 4.2. Third, we analyze the runtime efficiency of each algorithm in Section 4.3, capturing both merging cost and hyperparameter tuning overhead. Together, these evaluations provide a complete picture of the trade-offs between utility, robustness and computational efficiency across merging methods.
@@ 4.1 Multi-Task Performance
One of the primary advantages of model merging is its ability to combine the strengths of multiple specialized models into a single, multi-task model. Therefore, we first evaluate the multi-task performance of the merged models produced by different algorithms. Given the varying difficulty levels across tasks, we report normalized performance [26] as the main evaluation metric. Specifically, normalized performance is computed as $\frac{1}{n}\sum_{i\in[n]}\text{perf}^{(i)}_{\text{merged}}/\text{perf}^{(i)}_{\text{finetuned}}$, where $\text{perf}^{(i)}_{\text{merged}}$ and $\text{perf}^{(i)}_{\text{finetuned}}$ denote the performance of the merged and specialized models on task $i$, respectively. This metric captures the proportion of finetuned performance retained by the merged model, with a value of 1 indicating that the merged model matches the performance of the task-specific finetuned models across all tasks. We report the multi-task performance in Figure 2, and summarize our observations as follows. Full numeric reesults are presented in Appendix D.
@@ 4.2 Retention of Base Model Knowledge
Pretrained language models encode extensive knowledge acquired from large-scale, diverse training corpora, allowing them to generalize across a wide range of tasks. However, post-training can induce catastrophic forgetting, where useful capabilities of the base models are lost [36]. An additional advantage of model merging is its potential to mitigate forgetting compared to continual finetuning [25, 22]. Therefore, it is important to evaluate the extent of forgetting introduced by different merging algorithms to ensure that merged models not only excel on the five in-domain tasks but also retain generalization capabilities on unrelated tasks. To assess forgetting, we evaluate performance on a diverse set of benchmarks where the base models are known to perform well: i) General knowledge: MMLU [23], ii) Reading comprehension: TriviaQA [30] and SQuADv2 [52], iii) Domain-specific question answering: CoQA [53] and PubMedQA [28], iv) Translation: WMT 2014 French to English translation [6]. We demonstrate the trade-off of multi-task performance and forgetting in Figure 3.
@@ 4.3 Runtime
Figure 5: Performance versus wall-clock time (upper left better). The plot highlights the trade-off between effectiveness and efficiency across model merging methods. Both versions of Localize-and-Stitch, RegMean, and Task Arithmetic achieve a favorable balance.
Due to varying hyperparameter-tuning and training demands, merging algorithms exhibit markedly different running time. Since computational efficiency is a key advantage of model merging over traditional multi-task learning, we measure and report wall-clock time when merging Llama-3.2-3B models (Figure 5). For each algorithm, we separately report the total runtime in two components: i) algorithm runtime: the time required to execute the merging procedure, and ii) validation runtime: the time spent tuning hyperparameters (e.g., scaling factors, sparsity levels) on validation data with grid search. Although validation cost is often overlooked in prior work, we find it can dominate total runtime in real-world use cases.
@@
Efficiency-effectiveness tradeoff. To evaluate the practical utility of merging algorithms, we plot performance versus wall-clock time in Figure 5, where the top-left region represents the most desirable trade-off (high performance, low cost). Both versions of Localize-and-Stitch, RegMean, and Task Arithmetic achieve a favorable balance between effectiveness and efficiency. These methods consistently deliver strong performance without excessive runtime overhead.
@@ 5 Related Works
We compare MergeBench with prior evaluations of model merging in Table 4. Existing evaluations often lack either model diversity, sufficient scale, or support for complex merging algorithms.
@@ Appendix A Merging Methods Details
DARE [79] performs random dropout on the task vectors based on the per-task binary mask $m_{i}$ drawn from the Bernouli distribution $m\sim\text{Bernouli}(p)$, where $p$ is a predefined dropout rate. To retain the original scale, the dropout task vectors are rescaled by $1/(1-p)$. Overall, the resulting sparse task vectors are $\tau_{i}=(1-m_{i})\odot\tau_{i}/(1-p)$, and then DARE proceeds with the task arithmetic procedure to combine task vectors.
@@ B.1 Training data details
We present the training data statistics in Table 5.
@@ C.1 Training details
For 2B and 3B models, we conduct experiments on NVIDIA RTX A6000 GPUs using a sequence length of 4096 and an initial learning rate of 1e-5. For 8B and 9B models, we use NVIDIA A100 GPUs with a sequence length of 3072 and an initial learning rate of 5e-5. Across all model sizes, we adopt the AdamW optimizer [38] with a cosine learning rate schedule, and set the global batch size to 128. For all tasks, we perform SFT for 2 epochs.
@@ C.2 Hyperparameter Tuning
Merging algorithms have different requirements of hyperparameter tuning. We detail them in Table 6 following the practice specified in the original papers. For RegMean, although the original paper does not explicitly require hyperparameter tuning, we find that the selection of the scaling factor for the non-diagonal items in the inner product matrices dramatically influences the performance, and using the suggested $\alpha=0.9$ often results in poor performance. Thus, we treat it as a hyperparameter and perform validation. For Consensus TA, each of the 5 tasks require tuning of the sparsity parameters, and subsequently, it requires tuning the scaling coefficients, totaling 35 runs. For Dataless Localize-and-Stitch, the original paper suggests a sparsity of $5\%\sim 10\%$. However, at larger model scales, we find that effective localization requires activating more than 10% of the parameters. This may be explained by the observation that larger models tend to distribute knowledge more broadly across their parameters [2, 7], making task-specific information less concentrated. Consistent with this observation, Poppi et al. [48] report that identifying safety-relevant regions in multilingual LLMs requires localizing up to 20% of the parameters. These findings suggest that the sparsity requirements for effective merging may scale with model size. Thus, we treat sparsity as a hyperparameter and search from 10% to 50%.
Appendix D Full Numeric Results
We report the full numeric results of the multi-task performance of each merging algorithm in Table 7 and generalization performance in Figure 3.
@@ Appendix D Full Numeric Results
Appendix E Limitations
Our work shares some limitations with existing model merging literature. Firstly, our evaluation is limited to merging models finetuned from the same initialization, while how to effectively combine models from different base models remains an open problem. Secondly, our analysis and insights are mainly drawn from empirical observations, and it remains unclear whether the arguments can be tested theoretically. In fact, the theoretical understanding of the effectiveness of model merging is still underexplored, only with the few attempts from linear mode connectivity [15], tangent task spaces [42] and task relationships [81, 33].

### v3

@@ 1 Introduction
To address the limitations of existing model merging evaluations, we introduce MergeBench, a scalable and comprehensive benchmark designed to rigorously assess merging performance, illustrated in Figure˜1. First, MergeBench improves model selection by adopting state-of-the-art, open-source language models as base models. Specifically, we include both pretrained and instruction-tuned versions of Llama-3.2-3B, Llama-3.1-8B [17], Gemma-2-2B, and Gemma-2-9B [64], resulting in a total of eight base models. Second, we construct a more challenging and representative task suite for evaluating merged models. Each base model is further finetuned on one of five carefully selected task categories, including instruction following, mathematics, multilingual understanding, coding and safety, to produce specialized models with minimal skill overlap [^1] All of the 40 specialized models are open-sourced at https://huggingface.co/MergeBench.. By standardizing the finetuning and evaluation procedures, MergeBench ensures a fair and reproducible platform for comparing model merging algorithms. In addition to multi-task performance, MergeBench evaluates retention of pretrained generalization through forgetting analysis and reports runtime efficiency, offering a comprehensive view of both utility and computational cost of existing merging algorithms.
@@ 2 Background
Localize-and-Stitch [22]
i) Localization: Train binary mask to identify the most relevant parameters for each task.
Dataless localization: When no data available, retain largest top-$k$ parameters in task vectors.
ii) Stitch: Only stitch the localized regions back onto the base model.
Methods. We evaluate and compare eight representative model merging methods, summarized in Table˜1. The development of model merging techniques begins with the study of how to effectively assign merging coefficients to the constituent models. Model Soup, Task Arithmetic, Fisher Merging and RegMean exemplify this approach. As the field has evolved, researchers have identified sparsity in task vectors as critical to reducing interference during merging. Since finetuning often results in redundant or noisy parameter changes [45], sparse merging techniques aim to suppress uninformative updates. These methods typically involve sparsification alongside merging coefficient tuning. More details about each method is included in Appendix˜A.
@@ 3 MergeBench
MergeBench provides a framework to evaluate model merging methods with three key designs: task coverage, model selection, and training and evaluation procedure. We define diverse task categories (Section˜3.1), build specialized models from open-source LLMs (Section˜3.2), and apply standardized training and evaluation strategies (Section˜3.3) to ensure fair and reproducible evaluation.
@@ 3.3 Data Construction
Training. Since the base models already go through the pretraining stage, our training primarily focuses on post-training. For most task categories, we apply supervised finetuning (SFT) to align the base models to domain-specific behaviors. To better reflect realistic scenarios where specialized models may be developed asynchronously using different methods, we adopt additional training strategies where appropriate. Specifically, for mathematics on the 8B and 9B models, we further apply Group Relative Policy Optimization (GRPO) [57] on top of SFT to enhance the models’ reasoning capabilities. A summary of the training data and methods is provided in Table˜2, with detailed data statistics in Section˜B.1 and training configurations available in Section˜C.1.
@@
Merging. To enable fair comparison across merging algorithms, we standardize the evaluation protocol by unifying their data requirements. Except for Model Soup, all other methods depend on auxiliary data, which we categorize into two types. First, some methods require additional training data to compute model-specific statistics or perform optimization. This applies to Fisher Merging, RegMean and Localize-and-Stitch. For these methods, we uniformly sample 1,000 examples from the original training set. Second, several methods require validation data to tune hyperparameters, typically scaling factors that control the contribution of each task vector and sparsity levels that control the proportion of parameters retained. This category includes Task Arithmetic, TIES Merging, DARE, Consensus TA and Dataless Localize-and-Stitch. We tune these hyperparameters based on performance on surrogate validation tasks, with details in Section˜B.2.
@@
Evaluation. To assess the performance of the merged models, we curate a comprehensive evaluation suite covering all five task categories. The datasets and evaluation metrics used for each category are summarized in Table˜3. In addition to task performance, we evaluate the efficiency of each merging algorithm by reporting their wall-clock time, allowing for a holistic comparison that considers not only effectiveness but also the practical cost of applying each method.
@@ 4 Evaluation of Merging Methods
To provide a comprehensive evaluation of merging algorithms, we assess their performance along three key dimensions. First, we measure the multi-task performance of the merged models on the five target tasks, as detailed in Section˜4.1. Second, we assess forgetting of base model knowledge, evaluating how merging impacts the model’s generalization beyond the specialized tasks in Section˜4.2. Third, we analyze the runtime efficiency of each algorithm in Section˜4.3, capturing both merging cost and hyperparameter tuning overhead. Together, these evaluations provide a complete picture of the trade-offs between utility, robustness and computational efficiency across merging methods.
@@ 4.1 Multi-Task Performance
One of the primary advantages of model merging is its ability to combine the strengths of multiple specialized models into a single, multi-task model. Therefore, we first evaluate the multi-task performance of the merged models produced by different algorithms. Given the varying difficulty levels across tasks, we report normalized performance [26] as the main evaluation metric. Specifically, normalized performance is computed as $\frac{1}{n}\sum_{i\in[n]}\text{perf}^{(i)}_{\text{merged}}/\text{perf}^{(i)}_{\text{finetuned}}$, where $\text{perf}^{(i)}_{\text{merged}}$ and $\text{perf}^{(i)}_{\text{finetuned}}$ denote the performance of the merged and specialized models on task $i$, respectively. This metric captures the proportion of finetuned performance retained by the merged model, with a value of 1 indicating that the merged model matches the performance of the task-specific finetuned models across all tasks. We report the multi-task performance in Figure˜2, and summarize our observations as follows. Full numeric reesults are presented in Appendix˜E.
@@ 4.2 Retention of Base Model Knowledge
Pretrained language models encode extensive knowledge acquired from large-scale, diverse training corpora, allowing them to generalize across a wide range of tasks. However, post-training can induce catastrophic forgetting, where useful capabilities of the base models are lost [37]. An additional advantage of model merging is its potential to mitigate forgetting compared to continual finetuning [25, 22]. Therefore, it is important to evaluate the extent of forgetting introduced by different merging algorithms to ensure that merged models not only excel on the five in-domain tasks but also retain generalization capabilities on unrelated tasks. To assess forgetting, we evaluate performance on a diverse set of benchmarks where the base models are known to perform well: i) General knowledge: MMLU [23], ii) Reading comprehension: TriviaQA [30] and SQuADv2 [53], iii) Domain-specific question answering: CoQA [54] and PubMedQA [28], iv) Translation: WMT 2014 French to English translation [6]. We demonstrate the trade-off of multi-task performance and forgetting in Figure˜3.
@@ 4.3 Runtime
Figure 5: Performance versus wall-clock time (upper left better). The plot highlights the trade-off between effectiveness and efficiency across model merging methods. Both versions of Localize-and-Stitch, RegMean, and Task Arithmetic achieve a favorable balance relative to other methods.
Due to varying hyperparameter‑tuning and training demands, merging algorithms exhibit markedly different running time. Since computational efficiency is a key advantage of model merging over traditional multi-task learning, we measure and report wall-clock time when merging Llama-3.2-3B models (Figure˜5). For each algorithm, we separately report the total runtime in two components: i) algorithm runtime: the time required to execute the merging procedure, and ii) validation runtime: the time spent tuning hyperparameters (e.g., scaling factors, sparsity levels) on validation data with grid search. Although validation cost is often overlooked in prior work, we find it can dominate total runtime in real-world use cases.
@@
Efficiency-effectiveness tradeoff. To evaluate the practical utility of merging algorithms, we plot performance versus wall-clock time in Figure˜5, where the top-left region represents the most desirable trade-off (high performance, low cost). Both versions of Localize-and-Stitch, RegMean, and Task Arithmetic achieve a favorable balance between effectiveness and efficiency. These methods consistently deliver strong performance without excessive runtime overhead.
@@ 5 Related Works
We compare MergeBench with prior evaluations of model merging in Table˜4. Existing evaluations often lack either model diversity, sufficient scale, or support for complex merging algorithms.
@@ Appendix A Merging Methods Details
DARE [83] performs random dropout on the task vectors based on the per-task binary mask $m_{i}$ drawn from the Bernouli distribution $m\sim\text{Bernouli}(p)$, where $p$ is a predefined dropout rate. To retain the original scale, the dropout task vectors are rescaled by 1/(1p)1/(1-p). Overall, the resulting sparse task vectors are $\tau_{i}=(1-m_{i})\odot\tau_{i}/(1-p)$, and then DARE proceeds with the task arithmetic procedure to combine task vectors.
@@ B.1 Training data details
We present the training data statistics in Table˜5.
@@ C.1 Training details
Table 6: Comparison of performance across tasks for different model sizes.
Model Size Instruction Following Math Multilingual Coding Safety
2–3B models 10 40 5 18 24
7–8B models 36 154 17 66 82
For 2B and 3B models, we conduct experiments on NVIDIA RTX A6000 GPUs using a sequence length of 4096 and an initial learning rate of 1e-5. For 8B and 9B models, we use NVIDIA A100 GPUs with a sequence length of 3072 and an initial learning rate of 5e-5. Across all model sizes, we adopt the AdamW optimizer [39] with a cosine learning rate schedule, and set the global batch size to 128. For all tasks, we perform SFT for 2 epochs. We report the training cost in GPU hours for each task and model size in Table˜6.
@@ C.2 Hyperparameter Tuning
Merging algorithms have different requirements of hyperparameter tuning. We detail them in Table˜7 following the practice specified in the original papers. For RegMean, although the original paper does not explicitly require hyperparameter tuning, we find that the selection of the scaling factor for the non-diagonal items in the inner product matrices dramatically influences the performance, and using the suggested $\alpha=0.9$ often results in poor performance. Thus, we treat it as a hyperparameter and perform validation. For Consensus TA, each of the 5 tasks require tuning of the sparsity parameters, and subsequently, it requires tuning the scaling coefficients, totaling 35 runs. For Dataless Localize-and-Stitch, the original paper suggests a sparsity of $5\%\sim 10\%$. However, at larger model scales, we find that effective localization requires activating more than 10% of the parameters. This may be explained by the observation that larger models tend to distribute knowledge more broadly across their parameters [2, 7], making task-specific information less concentrated. Consistent with this observation, Poppi et al. [49] report that identifying safety-relevant regions in multilingual LLMs requires localizing up to 20% of the parameters. These findings suggest that the sparsity requirements for effective merging may scale with model size. Thus, we treat sparsity as a hyperparameter and search from 10% to 50%.
Appendix D Discussions and Limitations
Here we discuss about the scope of this project and limitations of existing model merging literature.
Firstly, our evaluation is limited to merging models finetuned from the same initialization. Merging models from different base models is beyond the scope of our work, as it introduces fundamentally different challenges from the model merging setting we study. In our definition, model merging refers to the technique that uses arithmetic operations in the model parameter space to combine the strengths of multiple models. This formulation is widely adopted in the literature, and is highly valuable in practical scenarios where training data is inaccessible and multiple teams finetune the same model in parallel. In contrast, merging across model families requires tackling architectural and tokenization mismatches, where parameter-level arithmetic is not directly applicable. Prior works have explored combining knowledge from heterogeneous models, such as model routing that dynamically selects among models at inference time [77], or model ensembling that aggregates outputs [81]. While important in their own right, these directions constitute separate problem formulations that are not directly comparable to our work. While extending merging to heterogeneous base models is a promising direction, we believe the within-family merging problem remains a rich and impactful domain with immediate utility.
Secondly, we focus on merging dense models, without considering Mixture-of-Experts (MoE) models. Merging MoE models introduces challenges fundamentally different from dense model merging due to their sparse activation pattern. A key issue is expert index mismatch: different MoE models may assign distinct meanings to the same expert index, and merging without alignment disrupts the routing semantics. As a result, merging of experts or router weights can misroute inputs to inappropriate experts, leading to degraded performance. Thus, dense-model merging techniques are not readily applicable to MoEs. Instead, existing approaches construct new MoE architectures by combining dense experts and routing among them [35, 60], which lies beyond our definition of model merging.
Thirdly, our analysis and insights are mainly drawn from empirical observations, and it remains unclear whether the arguments can be tested theoretically. Theoretical analysis of model merging is particularly challenging due to the scale and nonlinearity of modern transformer models, as well as the reliance on task-specific hyperparameter tuning. As a result, most progress in this area has been empirical, including our MergeBench, which is designed to systematically evaluate merging methods at scale. That said, we note that recent theoretical work has begun to analyze core components of model merging. For example, Li et al. [33] provides provable generalization guarantees for task vectors, showing that both low-rank approximation and magnitude-based pruning preserve performance, and that carefully chosen merging coefficients lead to strong generalization. These results support our empirical findings on the effectiveness of sparsification and coefficient tuning. only with the few attempts from linear mode connectivity [15], tangent task spaces [43] and task relationships [85, 33].
Appendix E Full Numeric Results
Overall performance. We report the full numeric results of the multi-task performance of each merging algorithm in Table˜8 and generalization performance in Figure˜3. As shown in Table˜8, merging Gemma models often yields stronger results than merging Llama models of similar size. While the precise cause remains unclear due to limited transparency into pretraining procedures, this suggests that certain model architectures or training pipelines may be inherently more merging friendly. This is an interesting direction for further investigation.
@@ Appendix E Full Numeric Results
Detailed per-task performance. To facilitate reproducibility of our evaluation results, we further report detailed per-task performance for all 8 models we test.
Table 10: Gemma-2-2B per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 19.6 29.4 22.6 24.6 19.8 26.3 26.3 17.0 23.1
Math 25.2 28.2 30.3 27.9 26.3 26.9 27.6 37.9 37.1
Multilingual 47.9 47.9 41.1 48.2 48.2 48.3 48.3 47.5 47.4
Coding 30.3 35.2 28.0 31.4 30.4 33.3 33.3 33.5 33.1
Safety 52.4 45.1 58.8 56.8 38.4 39.8 61.9 65.0 61.9
Avg. Acc 35.1 37.2 36.2 40.6 32.6 34.9 39.5 40.2 40.6
Avg. Norm 66.4 70.3 68.4 76.8 61.7 66.1 74.7 76.1 76.8
Table 11: Gemma-2-2B-IT per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 51.9 51.9 53.0 54.2 49.2 46.4 54.9 48.8 49.2
Math 38.7 38.7 39.7 40.5 38.5 36.7 38.4 47.9 47.9
Multilingual 49.2 49.2 48.7 48.7 49.3 49.0 49.2 48.8 48.7
Coding 40.2 40.2 40.1 39.3 39.6 38.3 39.7 40.2 38.2
Safety 81.3 81.3 76.8 83.6 76.3 74.8 81.0 79.8 79.0
Avg. Acc 52.3 52.3 51.7 53.3 50.6 49.0 52.7 52.7 52.6
Avg. Norm 89.9 89.9 88.8 91.6 87.0 84.3 90.5 90.6 90.4
Table 12: Llama-3.2-3B per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 7.2 25.3 12.0 14.2 9.6 18.7 30.5 10.4 23.7
Math 16.2 27.7 26.6 33.8 26.6 27.1 27.6 41.1 38.9
Multilingual 46.8 47.0 47.6 48.4 47.6 47.5 46.9 46.7 47.0
Coding 37.0 41.1 37.2 39.3 37.6 40.2 40.7 40.0 40.7
Safety 39.2 46.1 35.4 71.7 40.4 44.9 48.2 37.3 41.3
Avg. Acc 29.3 37.5 31.8 41.5 32.3 35.7 38.8 35.1 38.3
Avg. Norm 57.0 72.9 61.8 80.7 63.0 69.4 75.5 68.3 74.6
Table 13: Llama-3.2-3B-Instruct per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 56.0 59.7 53.6 56.9 56.6 48.6 58.8 55.8 57.2
Math 53.9 55.1 49.8 61.0 56.7 56.7 51.1 59.8 59.9
Multilingual 45.0 45.2 44.7 48.4 44.6 43.9 45.1 44.1 44.2
Coding 52.4 49.8 51.2 51.8 52.5 53.2 48.0 51.8 52.8
Safety 84.6 80.6 85.3 87.5 94.5 94.0 80.0 85.1 83.1
Avg. Acc 58.4 58.1 56.9 60.3 60.9 59.3 56.6 59.3 59.5
Avg. Norm 94.4 93.9 92.1 97.6 98.6 95.9 91.6 95.9 96.1
Table 14: Gemma-2-9B per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 30.3 31.2 27.5 33.8 28.8 32.0 23.7 30.5 35.3
Math 60.3 64.5 48.2 59.9 65.3 62.8 59.0 67.2 66.5
Multilingual 60.0 57.1 58.8 55.2 59.5 56.5 59.5 56.5 55.1
Coding 51.5 50.8 56.6 39.2 52.3 48.8 51.4 56.0 53.3
Safety 70.6 74.4 81.7 84.4 75.3 73.2 72.6 68.1 72.6
Avg. Acc 54.6 55.6 54.5 54.5 56.2 54.7 53.2 55.7 56.6
Avg. Norm 89.4 91.1 89.3 89.3 92.1 89.5 87.2 91.2 92.6
Table 15: Gemma-2-9B-IT per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 50.5 59.3 54.2 47.5 52.9 44.2 62.5 60.6 57.3
Math 64.4 64.3 55.5 52.4 66.3 62.5 63.8 70.4 67.7
Multilingual 60.9 63.0 58.6 49.5 60.6 57.6 63.3 63.3 57.7
Coding 58.5 59.8 58.9 48.8 59.5 55.7 60.5 59.4 55.6
Safety 68.2 75.3 74.7 73.1 71.6 62.4 77.3 79.0 69.7
Avg. Acc 60.5 64.4 60.4 54.2 62.2 56.5 65.5 64.6 61.6
Avg. Norm 98.3 104.6 98.1 88.2 101.0 91.8 106.5 104.9 100.1
Table 16: Llama-3.1-8B per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 8.3 31.2 5.2 10.9 12.2 13.1 26.3 18.1 37.3
Math 50.1 55.5 48.0 52.9 56.3 55.6 54.2 59.5 57.0
Multilingual 54.0 49.1 52.1 51.9 54.5 52.7 49.0 52.0 54.3
Coding 49.6 48.8 49.5 47.9 49.0 49.3 49.6 51.3 50.9
Safety 71.0 59.0 76.0 67.8 61.9 55.4 60.7 79.3 63.7
Avg. Acc 46.6 48.7 46.2 46.3 46.8 45.2 48.0 52.0 52.7
Avg. Norm 81.1 84.8 80.3 80.6 81.4 78.7 83.5 90.6 91.7
Table 17: Llama-3.1-8B-Instruct per-task and average results. Values are percentages.
Task Model soup Task arithmetic Fisher Merging RegMean TIES DARE Consensus TA Dataless L&S L&S
Instruction following 37.5 47.0 31.8 46.8 43.4 39.8 48.4 55.4 44.4
Math 64.4 60.3 52.3 59.9 65.7 63.5 63.3 68.2 67.1
Multilingual 53.6 54.8 54.3 50.9 53.9 51.4 54.7 54.2 52.3
Coding 62.1 61.8 63.6 58.0 62.6 57.3 61.2 62.9 62.0
Safety 81.4 79.8 86.2 89.4 90.4 87.8 79.6 80.4 91.8
Avg. Acc 59.8 60.7 57.6 61.0 63.2 59.9 61.4 64.2 63.5
Avg. Norm 88.6 90.0 85.4 90.4 93.7 88.8 91.1 95.2 94.1